OpenStudy (anonymous):

Let f(x)=sqrt(3x-6) Find f(14+h)= Find f(14)= Find f(14+h)-f(14)= Find (f(14+h)-f(14))/h= Use algebra to simplify the expression so that you can take the limit as . Right now if take that limit then you would wind up with the indeterminant form . Hint: h should only appear once in your simplified expression. Find f'(14)=(f(14+h)-f(14))/h as the limit of h approaches 0.

OpenStudy (anonymous):

So far I have: f(14+h)=sqrt(36+3h) f(14)=6 f(14+h)-f(14)=sqrt(36+3h)-sqrt(36) But I can't figure out the last two.

OpenStudy (anonymous):

f(14 + h) is simply the function with x replaced with 14 + h so the last two would be f(14+h)-f(14)= sqrt(3(14+h)-6) - sqrt(3(14)-6) ((f(14+h)-f(14))/h= ( sqrt(3(14+h)-6) - sqrt(3(14)-6) )/h this last one is important as it is the definition of a derivative I recommend memorizing it!

OpenStudy (anonymous):

A simple example f(x) = 3x thus f(x+b+c+d) = 3(x+b+c+d)

OpenStudy (anonymous):

hope this helps

OpenStudy (anonymous):

this is simple buddy

OpenStudy (anonymous):

all you have to do is multiply the numerator with its conjugate

OpenStudy (anonymous):

so the equation reduces to \[[ (36+3h)-36 ]\div(h (\sqrt{36+h}+6)\]

OpenStudy (anonymous):

now you can reduce the equation to \[3h \div(h \times(\sqrt{36+h} +6))\]

OpenStudy (anonymous):

h from numerator & denominator cancels. now you can directly apply the limits(put h=0) to find out the derivative

OpenStudy (anonymous):

also try to memorize the last equation as it is the first definition of derivative

OpenStudy (anonymous):

Most of the Mathematica attachment answers your questions.